# What Features in Prompts Jailbreak LLMs? Investigating the Mechanisms Behind Attacks

### v2

@@ 4.2 Non-linear probes achieve improved cross-layer generalization
Figure 4: Non-linear MLP probes (Bottom) transfer better to unseen layers than linear probes (Top). Transfer to other layers is measured by the total amount of probes that achieve $>80\%$ accuracy when predicting unseen layers (excluding the diagonal).
In this section, we investigate how probes trained on individual layers generalize when tested on other layers’ representations. Figure 4 shows cross-layer generalization matrices where each entry (i,j) represents the accuracy of a probe trained on layer i when tested on layer j. Several key patterns emerge:
First, both linear and non-linear probes show strong generalization to nearby layers, with accuracy degrading smoothly as the layer distance increases. This suggests that jailbreak-related features evolve gradually through the network rather than appearing abruptly at specific layers.
Secondly, non-linear probes achieve improved cross-layer generalization compared to linear probes. The diagonal patterns visible in the heatmaps demonstrate strong layer-to-layer correlations, but the spread of high-intensity values (shown in brighter colors) beyond the diagonal suggests that non-linear probes capture more complex relationships between distant layers.
This is particularly evident in the upper triangular regions of the matrices, where non-linear probes achieve notably higher accuracies compared to their linear counterparts. This asymmetric generalization pattern stems from probes generalizing better to later layers than to earlier ones. A probe trained on layer $i$ achieves higher accuracy when tested on layers $j>i$ compared to layers $j<i$. This progressive refinement of jailbreak features aligns with our finding that probe accuracy generally increases with layer depth (see Section 4.1).
@@ 4.3 Jailbreaking features are not universal: transfer to held-out attack methods is inconsistent
In this section, we investigate how well probes trained on all attack types perform on unseen / out-of-distribution (OOD) attack types. For this, we first filtered our dataset to only include the 10 most successful attacks. From here we then create ten different datasets, each excluding one attack type. We then split each attack dataset into train and test sets. The blue bars in Figure 5 represent performance on the test set including nine out of ten attack types. The red bars represent performance on only the held-out dataset, which includes only the unseen attack type. Out-of-distribution performance represented in red often falls behind the random guessing baseline. The mean accuracy of all probes (the mean of all red bars per figure) tends to be below $<$ 70 %, for all models except the not safety finetuned Ministral model, where the mean accuracy is clearly $>$ 70 % for all probe types. Overall, while probes sometimes do transfer successfully, they tend to not do so consistently for most models. This shows that successful jailbreaks from different methods attack the model using different, non-linear prompt features. However, there are clear differences in OOD performance of probes for different model architectures. In the next section, we show through causal intervention experiments that while non-linear probes do not tend to capture universal jailbreak features, they do seem to capture causally relevant ones.
@@ 4.4 Non-linear probes can guide effective mechanistic attacks and defenses
Finally, we use probes to design latent space perturbations (e.g., (Sheshadri et al., 2024; Arditi et al., 2024)) for the models to test whether the features identified by our probes correspond to causal mechanisms. We are interested in both offensive as well as defensive perturbations, increasing and decreasing compliance to harmful requests, respectively. We target all token positions at every step of generation in the model’s middle layer for all experiments.
We tune the strength of all perturbations via a hyperparameter grid search. We optimize interventions under the constraint that we maintain the general capabilities of the perturbed model, while ensuring minimal performance degradation as measured on the MMLU test set, a widely used capability benchmark (Hendrycks et al., 2021) (see Appendix E). Further, we performed a coherency check to assess answer quality using spaCy, evaluating multiple linguistic criteria including sentence structure, syntactic complexity, verb-noun distribution, and subject-verb agreement. We retained only hyperparameter configurations that produced $>90\%$ coherent answers on top of the retained MMLU performance. We evaluate the success of jailbreak attacks using the HarmBench autograder (Mazeika et al., 2024). The autograder considers a response a jailbreak if it does not refuse the request and produces harmful output aligned with the prompt. Incoherent or vague responses do not count, and the presence of caveats or errors does not disqualify an output if the requested behavior is still exhibited (see Appendix G for the specific prompt used for the autograder).
Our primary experiments in Figure 6, were conducted with the multilayer probes, that are trained on aggregated data across all layers. Single-layer probes, turned out to be less robust Appendix F.
Linear probe-guided interventions:
First, we investigate a linear probe-guided intervention. We add the weight vector of our linear probe to a given token position and layer of our model. Given a hidden layer $H$, we modify each token position $i$ corresponding to the prompt and generated response:
$h_{i}^{{}^{\prime}}=h_{i}\pm(w*\alpha)$
where $h_{i}$ is the original activation for token $i$ at layer $H$, $w$ is the coefficient vector of the linear classifier, and $\alpha$ is the intervention strength. This shifts each latent representation $h_{i}$ towards or against the direction that the linear probe associates with successful jailbreaks. Steering towards the jailbreak direction leads to offensive perturbations, and steering away from it leads to defensive perturbations.
@@
Figure 6: MLP probe-guided latent perturbations are more effective than linear probe-guided ones for controlling a model’s vulnerability to complying with harmful requests. Compared to any other method tested, MLP-guided offensive perturbations were the best attacks, and MLP-guided defensive perturbations were the best defenses. We restrict the strength of all latent perturbations to maintain approximate MMLU performance. In this figure, we present results for multilayer probes.
MLP and Transformer probe-guided interventions:
We extend the linear probe-guided latent-space attack to work with non-linear probes by iteratively performing gradient descent on a latent-space perturbation (see Figure 2 in Section 1 for an example attack). We iteratively perturb the hidden states $h_{1...n}$ so that the probe would classify them as successful jailbreaks for attacks or unsuccessful for defenses. Given a hidden layer $H$, we modify the activations at each token position $i$ using the gradient of the loss of the probe with respect to the model’s hidden states $h_{1...n}$. Given the loss, we compute the gradient with respect to the hidden layer activations and update the hidden activations using gradient descent:
$h_{i}{{}^{\prime}}=h_{i}-\eta\frac{\partial Loss}{\partial h_{i}}$
where $\eta$ is the learning rate. We optimize these perturbations over 32 steps and add them consistently to both prompt and response tokens of the model’s middle layer at each generation step.
After experimenting with multiple loss functions, we find the best results from a loss that moves the probe’s prediction a limited amount in the target direction. Specifically, we use mean squared error (MSE) loss to guide perturbations towards desired logit values.
@@
  • Offensive perturbations: Positive logit values (MLP probe predicts a successful jailbreak): we use an MSE loss toward a target of $\max(c_{o},p+c_{o})$ for some positive $c_{o}$.
  • Defensive perturbations: Negative logit values (MLP probe predicts an unsuccessful jailbreak): we use an MSE loss toward a target of $\min(-c_{d},p-c_{d})$ for some positive $c_{d}$.
We perform an automated grid search over different values of $c_{o}$, $c_{d}$, and learning rates for each model to find the configuration that leads to the most effective increase or decrease of compliance to harmful requests, while maintaining approximately equal capabilities as measured with the MMLU benchmark (see details in Appendix E). We further require generations to be coherent. To identify coherent datasets, we analyze the coherence of model outputs by applying a linguistic analysis pipeline that evaluates syntactic structure, sentence complexity, and subject-verb presence. Only datasets with at least 90% coherent outputs were included in the final set of results.
@@
We present results in Figure 6. Overall, MLP probe-guided perturbations were more effective than linear probe-guided perturbations and transformer probes. Offensive MLP probe-guided perturbations achieve higher attack success rates compared to linear probe-guided ones and the unperturbed baseline. Defensive perturbations of all probes achieved a lower compliance to harmful requests than the unperturbed baseline, and defensive MLP probe-guided perturbations achieve lower compliance to harmful requests compared to supervised safety finetuning. This indicates that the feature related to jailbreak success represented by the probes is not strictly linear. Notably, our interventions were also effective for Ministral-8B-Instruct, which shows that non-linear features related to jailbreaks exist even without explicit safety fine-tuning.
Non-linear probe-guided defenses outperform supervised safety fine-tuning: Finally, to test the practical competitiveness of non-linear-probe guided steering, we next compare it to the most common practical technique used to make systems robustly refuse harmful requests: fine-tuning. Here, we pit defensive linear, transformer, and MLP steering against supervised safety fine-tuning (SSFT) for all models. We finetuned models on the same dataset that the probes were trained on. In contrast to the probes, the model is trained not on latent representations but on prompts themselves. Labels are the ratings from the HarmBench autograder, indicating whether a jailbreak prompt lead to a jailbreak. We manually tuned the hyperparameters like learning rate and lora r, and trained for one epoch. Results in Figure 6 show that our MLP defense outperforms SSFT using the same data.
@@ 5 Discussion
Prompt embeddings are informative about vulnerabilities to jailbreaks. Our analysis sheds light on the mechanisms behind jailbreaking in large language models in several ways. First, our probing experiments demonstrate that jailbreak success can be predicted with high accuracy using both linear and non-linear probes trained solely on prompt representations. This suggests that signs to predict whether a model will be jailbroken can be found as early as in the model’s prompt encodings, even before the model generates a response.
Jailbreak-relevant features evolve gradually across layers. Our cross-layer analysis indicates that features related to jailbreak success are not static, but evolve dynamically through progressive transformations across the model’s layers. The asymmetric pattern of cross-layer generalization – where probes trained on later layers can effectively classify representations from earlier layers but not vice versa – indicates that the model gradually refines jailbreak-relevant features as information propagates through the network.
Jailbreaks cannot be understood in terms of linear or simple universal prompt features alone.
Our analysis reveals that probes exhibit poor generalization to unseen attack types, with linear probes particularly struggling to represent causally related features that can steer model behavior without capability trade-offs. The probes’ weak performance on held-out attacks indicates that each attack type may engage the model in a unique way and be represented differently internally. These findings suggest that different jailbreaking methods operate through distinct non-linear features rather than universal or linear ones. This suggests limitations with existing linear steering methods (e.g., (Arditi et al., 2024; Rimsky et al., 2023; Ball et al., 2024; Li et al., 2024b)), as they may not provide robust protection against diverse attack strategies.
Nonlinear probe-guided latent perturbations can be used effectively to diagnose harmful capabilities and defend against harmful behaviors. We find that MLP probes can be used to guide the design of offensive and defensive perturbations. We show that these perturbations improve over both existing linear perturbation methods and fine-tuning. This demonstrates that while there are only small observable differences in probe accuracy between linear and nonlinear probes, our nonlinear probe-guided perturbations can be a competitive, data-efficient method for evaluating and modulating LLM robustness.
@@
Limitations and future work: First, while we have shown that jailbreak features are non-linear and not universal, characterizing the precise nature of these features remains an open challenge. Secondly, it remains an open question whether defense strategies can effectively generalize to novel jailbreaking techniques, highlighting the need for more adaptable safety interventions.
Impact Statement
@@
Our work provides a deeper mechanistic understanding of jailbreaks in LLMs, offering insights into how harmful prompts exploit model vulnerabilities. This has implications for AI safety and robustness. While our findings can inform better defenses against adversarial attacks, they also highlight risks of exploitation if misused. Ethical considerations are crucial, as probing vulnerabilities must balance transparency with responsible disclosure. Given the limitations of current defenses in generalizing to unseen jailbreaks, our research underscores the need for adaptive security measures and regulatory oversight. Future work should focus on developing generalizable defenses and fostering collaboration between researchers, industry, and policymakers to ensure the responsible and ethical deployment of AI.
Appendix A Model Information
@@
Appendix B Attack Success Rates of Training Data
Appendix C Probe accuracies and Confusion Matrices of Multilayer Aggregated Probes
Figure 9: Multilayer probes predict jailbreak success with high accuracy based on prompt tokens. (see Figure 3 for single layer probes.)
Appendix D Illustration of MLP interventions
@@
Appendix E Capability
To assess their capabilities, all models with probe-induced activation modifications (shown in Figure 6) were evaluated on the MMLU benchmark. Models were retained only if their performance closely matched that of the unmodified baseline.
Figure 11: MMLU performance of model with probe interventions. Intervention hyperparameters were chosen in such a way to minimize capability tradeoffs as measured by the MMLU benchmark.
Appendix F Single Layer Probes

### v3

@@ 4.2 Non-linear probes achieve improved cross-layer generalization
We investigate how probes trained on individual layers generalize when tested on other layers’ representations. Figure˜4 shows cross-layer generalization matrices where each entry (i,j) represents the accuracy of a probe trained on layer i when tested on layer j. Several key patterns emerge: First, probes trained on layer $i$ perform better on deeper layers ($j>i$) than on earlier ones ($j<i$). Overall, accuracy increases with depth (Figures˜3 and 4; see Section˜4.1). Secondly, both linear and non-linear probes generalize well to nearby layers, with accuracy declining smoothly as layer distance increases. Models with lower multilayer probe accuracy (Gemma 7B Instruct and Qwen 2.5 7B Instruct) also showed weaker transfer in single-layer probes. Finally, non-linear probes generalize better across layers than linear ones. The broader spread of high-intensity values ($\star$) beyond the diagonal in Figure˜5 for MLP probes suggests they capture more complex, distributed features.
@@ 4.3 Jailbreaking features are not universal: transfer to held-out attack methods is inconsistent
We evaluate how well probes trained on known attacks generalize to unseen (OOD) attack types. Using the 10 most successful attacks, we create ten datasets, each holding out one attack. Each dataset is split 80/20 into train and test sets.
We evaluate probe generalization to unseen attack types by holding out one attack at a time Figure˜5 . Across models, probe accuracy on out-of-distribution (OOD) attacks often falls below the random guessing baseline, with mean accuracy typically under 70% and only a small portion of probes reaching a high accuracy of $80\%$. An exception is the non safety-finetuned Mistral model, where all probe types achieve consistently higher OOD accuracy. These results suggest that different jailbreak methods exploit distinct, non-linear prompt features, and that probe generalization varies substantially across model architectures. In the next section, we show through causal intervention experiments that while non-linear probes do not tend to capture universal jailbreak features, they do seem to capture causally relevant ones.
@@ 4.4 Non-linear probes can guide effective mechanistic attacks and defenses
Figure 6: MLP probe-guided latent perturbations are more effective than linear probe-guided ones for controlling a model’s vulnerability to complying with harmful requests. Compared to any other method tested, MLP-guided offensive perturbations were the best attacks, and MLP-guided defensive perturbations were the best defenses. We restrict the strength of all latent perturbations to maintain approximate MMLU performance (see Section˜F.2. In this figure, we present results for multilayer probes. See Appendix˜H and Appendix˜I for further experiments.
Finally, we use probes to design latent space perturbations (Sheshadri et al., ; Arditi et al., ; Braun et al., ) to test whether the identified features are causally linked to model behavior. We apply both offensive and defensive perturbations, respectively increasing or decreasing harmful compliance. Perturbations target all token positions at each generation step in the model’s middle layer. Our experiments in Figure˜6, were conducted with the multilayer probes, that are trained on aggregated data across all layers. Single-layer probes, turned out to be less robust Appendix˜I.
To ensure that interventions preserve model quality, we perform a hyperparameter grid search over perturbation scales $\alpha\in\{-1,-0.5,-0.4,-0.25,-0.2,0.2,0.25,0.4,\\ 0.5,1\}$ for linear probes, and over target values $t\in\{40,30,20,15,10,5\}$ for nonlinear (MLP and Transformer) probes. For the nonlinear probes, we also tune the learning rate $\eta\in\{0.01,0.025,0.05,0.1\}$. In linear probes, negative $\alpha$ values correspond to defensive perturbations and positive values to offensive ones. For MLPs, each target value is tested in both directions. We then select configurations that minimize performance degradation on the MMLU benchmark (Hendrycks et al., ) (see Section˜F.2). Jailbreak success is evaluated using the same HarmBench autograder (Mazeika et al., ) used to label our training data. The autograder flags outputs as jailbreaks if they fulfill the harmful request, even when caveated, but excludes incoherent or vague completions (see Appendix˜J).
Linear probe-guided interventions.
We first investigate interventions guided by a linear probe. At a selected layer $H$, we modify the representation of each token position $i$ (from both the prompt and generated response) as $h_{i}^{\prime}=h_{i}\pm\alpha w$, where $h_{i}$ is the original activation at position $i$, $w$ is the weight vector of the trained linear classifier, and $\alpha$ is the perturbation strength. This shifts each latent representation either toward or away from the direction associated with jailbreak success: positive $\alpha$ values apply offensive perturbations (increasing jailbreak likelihood), while negative values apply defensive perturbations (reducing it).
MLP and Transformer probe-guided interventions.
We extend the linear probe-guided attack to nonlinear probes by using gradient descent to iteratively adjust latent activations in the direction the probe associates with jailbreak success or failure (see Figure˜2 in Section˜1 for an example attack). Specifically, we perturb the hidden states $h_{1},\dots,h_{n}$ so that, if these modified representations were passed to the probe, it would classify them as either a successful jailbreak (for offensive interventions) or an unsuccessful one (for defenses). At a selected hidden layer $H$, we compute the gradient of the probe’s loss with respect to each token’s hidden activation and apply a gradient descent update: $h_{i}^{\prime}=h_{i}-\eta\,\frac{\partial\mathcal{L}}{\partial h_{i}}$, where $\eta$ is the learning rate and $\mathcal{L}$ is the probe’s classification loss. This update is repeated multiple times, and the resulting perturbations are applied to all prompt and response tokens at each generation step in the model’s middle layer. After experimenting with multiple loss functions, we find the best results from a loss that moves the probe’s prediction a limited amount in the target direction. Specifically, we use mean squared error (MSE) loss to guide perturbations towards desired logit values.
In summary, we guide nonlinear interventions using an MSE loss on the MLP probe’s logit output $p$, pushing it toward a target value.
@@
Offensive: Push $p$ toward $\max(t_{o},p+t_{o})$, with positive target $t_{o}$, encouraging the probe to predict a successful jailbreak.
Defensive: Push $p$ toward $\min(t_{d},p-t_{d})$, with negative target $t_{d}$, encouraging the probe to predict a failed jailbreak.
@@
Overall, MLP probe-guided perturbations were more effective than linear and transformer-based ones (see Figure˜6). Offensive MLP perturbations achieved higher attack success rates than both linear perturbations and the unperturbed baseline. Defensive perturbations reduced harmful compliance across all probe types, with MLP-guided defenses proving most effective. To contextualize these results, we also compared MLP, transformer, and linear defenses to a simple supervised safety fine-tuning (SSFT) baseline. While improving jailbreak resistance is not the focus of this work, we finetuned models on the same data used to train the probes, using HarmBench labels and a single training epoch. Notably, defensive MLP perturbations outperformed SSFT, indicating that our method can be considered effective relative to this baseline. Finally, our methods remained effective even on Mistral-8B-Instruct, which lacks explicit safety fine-tuning, indicating that nonlinear jailbreak-relevant features are present even in models without alignment training.
@@ 5 Discussion
Prompt representations encode early jailbreak signals.
Our results (Figure˜3) show that prompt representations – before any output is generated – contain sufficient information to predict jailbreak success. Probes trained solely on these prompt-side latents classify jailbreak success with high accuracy, suggesting that LLMs internally encode signals of vulnerability upstream of any response. Moreover, as shown in Figure˜5, these features evolve across depth: cross-layer generalization is asymmetric, with probes trained on earlier layers transferring better to later ones, and deeper layers encoding more robust jailbreak signals overall.
Jailbreak-relevant features are model-specific and non-linear.
Our cross-layer results Figure˜4 show that non-linear probes, particularly MLPs, generalize more effectively to nearby and deeper layers than linear probes, suggesting they capture richer, more distributed features. Still, probe generalization degrades with layer distance, and performance on unseen attack types is consistently low. This indicates that jailbreak success is mediated by complex, non-overlapping features specific to each attack type and model, rather than a shared or universal latent signature. Linear probes, in particular, struggle to capture features that are both predictive and causally effective for steering model behavior without degrading capabilities (Figure˜6), highlighting limitations of linear steering approaches (e.g., (Arditi et al., ; Rimsky et al., ; Ball et al., ; Li et al., 2024)). This asymmetry between attack and defense mechanisms is noteworthy: while Zheng et al. found that safety prompts operate through relatively coherent directional shifts that can be captured with low-dimensional linear methods, our results suggest that successful jailbreaks exploit more heterogeneous, attack-specific vulnerabilities that resist simple linear characterization – suggesting limitations to universal defense strategies that DRO implicitly assumes. Our findings challenge the assumption that jailbreak-related features are universal across attack types or easily isolated using simple linear methods (Zou et al., 2025; Ball et al., ; Arditi et al., ; Li et al., ; Li et al., ; Park et al., ; Bhattacharjee et al., ; Stickland et al., ). Instead, they point toward a more complex landscape of jailbreak vulnerabilities.
Non-linear probes enable causal insights, not just steering.
While differences in probe classification accuracy between linear and non-linear models are modest, the corresponding interventions differ substantially. MLP-guided perturbations are more effective at increasing or suppressing jailbreak success, without harming model capabilities. This suggests that jailbreak-relevant features are non-linear and distributed in latent space, and that non-linear probes can reliably uncover and modulate them. We frame our perturbation method not primarily as a defense mechanism, but as a tool for scientific analysis, enabling causal interrogation of internal representations. In this light, non-linear probes offer a promising approach for studying how harmful behaviors are encoded in LLMs, and for developing more interpretable safety interventions.
6 Limitations
While our findings suggest that different jailbreak methods rely on distinct, non-linear features, this does not rule out the existence of universal representations. It is possible that such features exist but are not captured by the probe architectures used here. Our use of “universal” is therefore practical: we refer to features simple enough to be detected by current probing methods, not to their theoretical non-existence.
@@ 6 Limitations
In addition, although we show that linear and universal explanations are insufficient, the precise nature of jailbreak-relevant features remains unclear. Finally, it remains an open question whether defense strategies can effectively generalize to future jailbreaking techniques, highlighting the need for more adaptive safety interventions.
Acknowledgments
This work was partially supported by the Cambridge ERA:AI Fellowship and the UK Research and Innovation [grant number EP/S023356/1] in the UKRI Centre for Doctoral Training in Safe and Trusted Artificial Intelligence (www.safeandtrustedai.org);
We are grateful to the Center for AI Safety for generously providing computational resources needed to run many of the experiments in this paper.
Appendix A Impact Statement
@@
Our work provides a deeper mechanistic understanding of jailbreaks in LLMs, offering insights into how harmful prompts exploit model vulnerabilities. This has important implications for AI safety and robustness. While our findings can inform the design of better defenses against adversarial attacks, they also expose potential risks if such insights are misused. Ethical considerations are crucial, as probing vulnerabilities must balance transparency with responsible disclosure. Given the limitations of current defenses in generalizing across attack types, our research underscores the need for more adaptive security measures and regulatory oversight. We advocate for future research that develops generalizable defenses while fostering collaboration across academia, industry, and policymakers. Ensuring the safe and ethical deployment of LLMs will require not only technical innovation but also coordinated oversight and a strong commitment to responsible AI governance.
Appendix B Model Information
@@
Appendix C Attack Success Rates of Training Data
(a) Gemma-7B-Instruct
(b) LLaMA-3.1-8B-Instruct
Figure 7: Attack Success Rates (ASRs) by jailbreak type across different models.
(a) LLaMA-3.2-3B-Instruct
(b) Mistral-8B-Instruct
(c) Qwen-2.5-7B-Instruct
Figure 8: Attack Success Rates (ASRs) by jailbreak type across different models.
Appendix D Probe accuracies and Confusion Matrices of Multilayer Aggregated Probes
(a) Gemma 7B: Linear Probe
(b) Gemma 7B: MLP Probe
(c) Gemma 7B: Transformer Probe
(d) LLaMA 3.1: Linear Probe
(e) LLaMA 3.1: MLP Probe
(f) LLaMA 3.1: Transformer Probe
(g) LLaMA 3.2: Linear Probe
(h) LLaMA 3.2: MLP Probe
(i) LLaMA 3.2: Transformer Probe
Figure 9: Multilayer probes predict jailbreak success with high accuracy based on prompt tokens. Each subfigure shows a confusion matrix for a probe type and model. See Figure˜3 for single-layer results.
(a) Mistral: Linear Probe
(b) Mistral: MLP Probe
(c) Mistral: Transformer Probe
(d) Qwen 2.5: Linear Probe
(e) Qwen 2.5: MLP Probe
(f) Qwen 2.5: Transformer Probe
Figure 10: Multilayer probes predict jailbreak success with high accuracy based on prompt tokens. Each subfigure shows a confusion matrix for a probe type and model. See Figure˜3 for single-layer results.
Appendix E Illustration of MLP interventions
@@
Appendix F Capability
F.1 Linguistic Coherence Filtering and Capability Preservation
To assess the linguistic coherence of model outputs, we applied a lightweight syntactic analysis pipeline using the spaCy library with the en_core_web_sm English model.
Each model response was processed as follows:
  • Tokenized and parsed into sentences using spaCy’s dependency parser.
  • Counted for core part-of-speech (POS) categories, including nouns and verbs.
  • Analyzed for syntactic structure, particularly the presence of subject-verb relationships, identified using dependency labels (e.g., nsubj and VERB/AUX POS tags).
We marked a response as coherent if it satisfied all of the following conditions:
  • Contains at least one sentence,
  • Has an average sentence length exceeding 3 tokens,
  • Contains at least one verb and one noun,
  • Includes at least one sentence with an explicit subject-verb dependency relationship.
Any dataset used in this paper was included only if it contained at least 90% coherent responses according to these criteria. For our main experiments (Figure˜6), we further required that interventions cause no meaningful reduction in model capability, as measured by performance on the MMLU benchmark.
F.2 MMLU Benchmark
For our main set of experiments (Figure˜6), we imposed an additional constraint: models were required to show no effective capability reduction on the MMLU benchmark after intervention. This ensured that we evaluated all models at a comparable level of general effectiveness. By filtering out interventions that significantly degraded MMLU performance, we isolated the effects of our methods on jailbreak susceptibility without conflating them with broader reductions in model utility.
To assess their capabilities, all models with probe-induced activation modifications (shown in Figure˜6) were evaluated on the MMLU benchmark. Models were retained only if their performance closely matched that of the unmodified baseline.]
Figure 12: MMLU performance of the model with probe interventions. Intervention hyperparameters were chosen to minimize capability tradeoffs, as measured by the MMLU benchmark.
Appendix G Autograder Areement
To assess the validity of our evaluation framework, we compared the ratings of the HarmBench (Mazeika et al., ) used throughout the paper and StrongReject (Souly et al., ), another popular autograder. HarmBench provides binary labels (Yes/No) indicating whether a model’s output constitutes a strong rejection of harmful content. In contrast, StrongReject assigns a continuous score between 0 and 1 reflecting the probability that a response constitutes a strong rejection.
For comparability, we thresholded StrongReject scores to produce binary labels, and computed agreement metrics between the two systems. We evaluated this agreement on responses from five prominent LLMs: LLaMA-3 7B, LLaMA-3 8B, Gemma 7B, Mistral 7B, and Qwen 7B. Each model’s outputs were passed to both autograders, and comparisons were performed at the dataset level.
About StrongReject.
StrongReject is an automated evaluator introduced by Mazeika et al. . It provides a continuous score based on whether a response contains harmful content, judging both refusal and content quality. The original implementation used GPT-4 Turbo and scored responses based on refusal, convincingness, and specificity. In our study, we utilized the fine-tuned Gemma 2B version of StrongReject to evaluate model responses.
Figure˜13 shows the relationship between HarmBench ASR and StrongReject ASR across all datasets and models. We observe a strong overall correlation ($\rho=0.82$), indicating that both autograders largely agree on which responses are classified as strong rejections. This high agreement also holds within individual models, with most points lying close to the diagonal. These results support the reliability of StrongReject as a proxy for HarmBench in our evaluations.
Figure 13: Dataset-level agreement between HarmBench and StrongReject autograders. Each point corresponds to a dataset-model pair. The diagonal represents perfect agreement. The high Spearman correlation ($\rho=0.82$) indicates substantial alignment between the two evaluation methods.
Appendix H Intervening on Different Layers
We conducted additional experiments by applying causal interventions using multilayer probes at various depths within the model: early, middle, and late layers. This analysis complements our main results by examining how the layer position influences the effectiveness of interventions.
(a) Early layer
(b) Middle layer (reproduced from Figure˜6)
(c) Late layer
Figure 14: Causal interventions using multilayer probes applied at early, middle, and late layers of the model. Across all depths, MLP-based defenses consistently achieve the highest compliance reduction, while MLP-based attacks are often more effective or comparable to linear probes. Unlike our main experiments, these interventions were evaluated only on answer coherency without enforcing the MMLU capability constraint discussed in Section F.2.
Appendix I Single Layer Probes